OpenStudy (mathmath333):

Mithilesh makes 750 articles at a cost of 60 paise per article.He fixes the selling price such that if only 600 articles are sold,he would have made a profit of 40% on the outlay.However 120 articles got spoilt and he was able to sell 630 articles at this price.Find his actual profit percent of total outlay assuming that he unsold articles are useless. a.42% b.53% c.47% d.46%

ganeshie8 (ganeshie8):

so he makes a profit of 40% of his total investment just by selling 600 articles

ganeshie8 (ganeshie8):

suppose selling price of each aticle = \(p\)

ganeshie8 (ganeshie8):

can we have \[600\times p = 750\times 60 \times 1.4 \] ?

ganeshie8 (ganeshie8):

solving that gives you selling pice of each article : \(p = 105\) paise rest should be easy..

OpenStudy (mathmath333):

u mean \(\dfrac{105-60}{60}\times 100=75\%~---~?\)

ganeshie8 (ganeshie8):

yes thats the projected profit if he sells all 750 articles

ganeshie8 (ganeshie8):

he sold only 630 articles so you need to be a bit careful

ganeshie8 (ganeshie8):

its simple algebra profit = (total sales revenue) - (total cost)

ganeshie8 (ganeshie8):

total sales revenue = 630*105 = ? total cost = 750*60 = ?

OpenStudy (mathmath333):

ok thanks